# arXiv:2602.13749v3[cond-mat.str-el] 02 Mar 2026

### v3

@@ Abstract
Non-Hermitian multichannel Kondo problems host both non-Fermi liquid and non-Hermitian physics, which provide a prototypical model to explore exotic collective quantum phenomena driven by the two different ingredients. Here, we first propose an experimental setup that realizes this model with exact channel symmetry as well as a controllable $\mathcal{PT}$ symmetry. Then, we perform a multi-method study of this model, focusing on the low-energy spectrum, the thermodynamic quantities, and the transport properties associated with different fixed points. Using the Bethe ansatz approach, we identify existence of the Yu-Shiba-Rusinov-like state previously found in the non-Hermitian single-channel Kondo model. Then, based on non-Hermitian numerical renormalization group calculations, we reveal clear numerical signatures of the Yu-Shiba-Rusinov state emerging in the relatively strong non-Hermiticity regime of the $\mathcal{PT}$-asymmetric model. Furthermore, our boundary conformal field theory, which is found to be applicable for the $\mathcal{PT}$-symmetric model, uncovers an anomalous temperature dependence of the Kondo conductance, which is beyond conventional Hermitian Kondo systems.
@@ I Introduction
Non-Hermitian quantum models, which physically arise from the loss or gain via coupling to an external environment, are well known to host a wealth of quantum phenomena absent in the Hermitian counterparts, e.g., the non Hermitian skin effect [1, 2, 3, 4, 5, 6, 7, 8, 9], non-Hermitian topology [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], exceptional points [20, 21, 22, 23, 24] and unexpected dynamics [25, 26, 27, 28, 29, 30, 31]. These findings open a new avenue in which many key concepts or tools for closed systems find new applications. However, most existing works investigate non-Hermitian effects in non-interacting or weakly interacting systems, where quasi-particle descriptions are well-defined and applicable. Although there has been some studies on correlated non-Hermitian systems, they mainly focus on specific solvable models [32, 33, 34, 35, 36, 37] as well as their ground state properties. The question that how the non-Hermiticity would affect correlated systems without well-defined particles has not be addressed yet. Particularly, it remains unclear whether or not any universal collective behavior would emerge in such cases.
@@
A typical and important example that violates the quasi-particle description is the non-Fermi liquid (NFL). It is one of the core concepts in modern condensed matter physics, describing a novel class of emergent phenomena beyond the Landau Fermi liquid paradigm. In strong-correlated systems where the conventional quasi-particle picture breaks down, there could emerge exotic collective behaviors, such as the unconventional scaling of thermodynamic quantities versus temperatures. A thorough understanding of these correlated phenomena are of key importance, as they are fundamentally related to the high-temperature superconductors [38, 39, 40, 41, 42, 43, 44], quantum spin liquids [45, 46, 47, 48, 49, 50], topological orders [51, 52, 53, 54], etc. On the other hand, the mechanism driving the non-Fermi liquid behavior can be various. It could arise from strong coupling between electrons and critical bosons, thus could naturally take place around quantum criticalities [55, 56, 57]. Besides, it can also be driven by the proximity to quantum spin liquids [58, 59, 60, 61, 62, 63], which indicates that the frustration and correlation are essential for its generation.
@@
Experimental signatures of NFL have been observed in various correlated systems, including heavy fermion materials [64, 65, 66, 67, 68, 69, 70, 71, 72, 73], high-$T_{c}$ superconductors [38, 39, 40, 41, 42, 43, 44], etc. Among them, the multichannel Kondo effect has attracted enormous interest, as it serves as prototypical platform for the experimental control over the NFL behaviors [74, 75, 76, 77, 78, 79]. To realize a multichannel Kondo system that drives the NFL fixed point, exact channel symmetry should be achieved. Although this has been experimentally demonstrated by fine-tuning in nanostructuress [74, 75], another remarkable proposal is to utilize the topological Kondo effect. Based on multi-junction setup where Majorana fermions tunnel into normal leads, Béri and Cooper demonstrates a feasible and controllable way to realize the multi-channel Kondo model with exact channel symmetry without the need of fine-tuning [80]. Considering that the multichannel Kondo model hosts non-Fermi liquid fixed point, which exhibits fractionalized degrees of freedom contributing to a fractional impurity entropy [81, 82, 83], it is intriguing and worthwhile to ask would any new collective physics be emerging when non-Hermiticity further comes into play?
@@
Despite the multichannel feature, the non-Hermitian single channel Kondo problem has been proposed and studied in previous works  [35, 37]. Ref. [37] carried out a perturbative renormalization group calculations and found that there occurs two fixed points, respectively in the weak- and strong-coupling regime. Due to the non-Hermitian effects, there occurs a circling RG flow around the weak-coupling fixed point, driving the system towards the local moment phase. Moreover, it is argued that the strong-coupling regime describes the Kondo phase, in analogy with conventional Kondo physics of the Hermitian counterpart. A more recent study [35] then performed a more careful solution of the Bethe ansatz solution. Remarkably, it reveals a more interesting phase, where the impurity is screened by effective extra degrees of freedom arising from the non-Hermitian effect, forming a single-mode bound state. Owing its similarity with the Yu-Shiba-Rusinov (YSR) impurity states in conventional $s$-wave superconductors, it is dubbed the YSR phase. These results indicate that the non-Hermiticity brings about a richer phase diagram than expected, which worth a detailed and comprehensive study by non-perturbative methods.
@@
In this work, we perform a comprehensive, multi-method investigation of the non-Hermitian Kondo model, with a particular focus on the multichannel cases. Our aim here is threefold. First, although the multi-junction setup [80] can be used to realize the conventional Hermitian multichannel Kondo model, it remains unclear how to systematically introduce the non-Hermitian effects into the system. Thus, we propose a realistic setup and demonstrate a step-by-step realization scheme. Second, we use a number of methods to thoroughly study the non-Hermitian multichannel Kondo (NHMCK) model in order to achieve a more comprehensive understanding beyond perturbative perspectives, involving the perturbative renormalization group (RG), Bethe ansatz, non-Hermitian numerical renormalization group (NHNRG), non-Hermitian Kubo formula, as well as the boundary conformal field theory (BCFT). Particularly, we clarify the role of $\mathcal{PT}$ symmetry by considering both the $\mathcal{PT}$-symmetric and $\mathcal{PT}$-asymmetric NHMCK. Different features in terms of the low-energy spectrum and the thermodynamic quantities are found for the two cases. For the $\mathcal{PT}$-asymmetric NHMCK, our NHNRG method reveals clear numerical evidence about the YSR phase predicted by Ref. [35]. For the $\mathcal{PT}$-symmetric model, we combine the NHNRG with an additional analytic proof and show that the low-energy physics in this case is well within the BCFT descriptions. Third, we go beyond previous work and explore the nontrivial transport behaviors generated by the NFL and the non-Hermitian physics. We find that both the weak-coupling and the strong-coupling fixed point exhibit unconventional Kondo conductance different from the Hermitian systems. Around the weak-coupling FP, the impurity moment decouples from the conduction electrons even for antiferromagnetic Kondo couplings, leading to a universal low-temperature conductance, $\sim 1/\mathrm{ln}^{2}(T/T_{K})$. Although the conductance is similar to that of the underscreened Kondo models [84, 85, 86], it is driven by a completely different dissipation mechanism–the nH-induced decoupling. Remarkably, around the strong-coupling fixed point, the BCFT analysis predicts an anomalous deviation in the Kondo conductance with raising the temperature. Such a non-Hermiticity enriched anomaly stems from an emergent BCFT with non-Hermitian boundary operators, which is unique to systems exhibiting both non-Hermiticity and non-Fermi liquid behaviors.
@@
Last, we mention that this present work is companion to another work by us, i.e., Ref. [87]. More concise results with more clear pictures are discussed in Ref. [87], while the technical aspects in terms of methods and more detailed discussions are included in this manuscript.
@@
The remaining part of this work is organized as follows. In Sec. II, we introduce the model Hamiltonian and demonstrate a step-by-step realization scheme of the NHMCK model. In Sec. III, we outline the results given by the perturbative RG calculations, which suggests two possible phases corresponding to the weak- and strong-coupling fixed points. In Sec. IV, we present the Bethe ansatz calculations that reveals existence of a YSR state, which is beyond the perturbative RG analysis in Sec. III. In Sec. V, we perform a NHNRG study of the model. We first introduce the technical details in the implementation of the NHNRG method and then utilize it to investigate the $\mathcal{PT}$-asymmetric NHMCK model. In Sec. VI, our focus is on the pseudo-Hermitian model with $\mathcal{PT}$ symmetry. We firstly show the NHNRG results about the low-energy spectrum and the thermodynamic quantities, which justify the validity of the strong-coupling Kondo phase predicted by the perturbative RG. Then, we outline a more rigorous proof showing that the BCFT approach still applies for this model with $\mathcal{PT}$ symmetry. In Sec. VII, we explore the transport properties associated with both fixed points. For the weak-coupling fixed point, we present the calculation procedures based on the non-Hermitian Kubo formula, which is well applicable for the perturbative regime. Then, for the strong-coupling fixed point, we exploit the BCFT, and demonstrate the anomalous Kondo conductance generated by the interplay between non-Hermiticity and non-Fermi liquid physics. More discussions and conclusions are included in Sec. VIII.
II Model Hamiltonian and possible realization setup
@@
Figure 1: The schematic plot of the realization setup of the NHMCK model. (a) The multi-junction nanostructure, where the Majorana modes (the red dot) on a superconducting island are connected to normal leads. The tunneling between the Majorana mode and the lead is assisted by a quantum do (the orange dot). (b) A zoom-in plot of the Majorana-lead tunneling junction via the quantum dot. The quantum dot is further coupled to an environment. (c) The quantum dot has a single-body loss into the environment and monitored by a reservoir, and a post-selection is implemented on the Lindbladian dynamics of the quantum dot.
The tunneling junction is plot in more details in Fig.1(b), which involves the dissipative QD. The latter consists of a QD coupled to environment. The QD-environment coupling system is further subjected to a monitoring-feedback device, which resets the particle loss events and ensures the particle conservation of the QD, as shown by Fig.1(c).
@@
where $\varrho$ is the reduced density matrix of the QD, $H^{\prime}=H_{\mathrm{QD}}+H_{\mathrm{LS}}$ where $H_{\mathrm{LS}}$ is the Lamb-shift term taking the form of $H_{\mathrm{LS}}=\sum_{i}\Omega_{i}d_{i}^{\dagger}d_{i}$, with $\Omega_{i}$ being a pesudo-potential for QD acquired from coupling to the environment. The jump operators $L_{i}$, which originate from the coupling of the quantum dot operators to the bath, can be constructed through the following manner, [92, 90, 93],
@@
In order to realize an non-Hermitian system, the particle leakage event should be monitored by an attached empty reservoir as a particle detector [89, 95] and then an immediate resetting needs to be implemented by the feedback equipment. An alternative option is to take the record in the time period when no particle leakage is monitored. These procedures result in a postselection on the Lindbladian, leading to an effective non-Hermitian Hamiltonian description of the QD, i.e.,
@@
where $\tilde{\epsilon}_{d,i}$ denotes the Lamb-shifted QD energy level. It is clear from above that the coupling to the envirnoment and the postselection generate an dissipative QD with an imaginary chemical potential. Note that the effective non-Hermitian Hamiltonian description can be also accessed through the Green’s function point of view (see Appendix A).
@@
This normalized density matrix will be used latter in our calculation of Kondo conductance, as will be shown in Sec. VII.
@@
where we have focued on the single-mode case with $N_{d}=1$ thus neglected the index $i$ for simplicity. The Majorana-lead tunneling is given by,
@@
We now show that the dissipative effect of the QD can induce an non-reciprocal tunneling process of the Majorana fermions. This can be read off from the Schrödinger equation of the junction system, i.e., $H_{\mathrm{junc}}|\Psi\rangle=E|\Psi\rangle$, which is written as a coupled equation in the direct product space of the QD and the rest part. Tracing out the QD degrees of freedom then generates an effective junction Hamiltonian renormalized by the dissipative QD as,
@@
After insertion of Eq.(8), we find that the original tunneling amplitude is renormalized in a nontrivial way. The tunneling from the Majorana to the lead ($t$) and the reversed tunneling $(t^{\star})$ are respectively modified as (see Appendix B),
@@
III Perturbative renormalization group results

### v4

# Interplay between non-Fermi liquid and non-Hermiticity: A multi-method study of non-Hermitian multichannel Kondo model

@@ Abstract
Non-Hermitian multichannel Kondo problems host both non-Fermi liquid and non-Hermitian physics, which provide a prototypical model to explore exotic collective quantum phenomena driven by the two different ingredients. Here, we first propose an experimental setup that realizes this model with exact channel symmetry as well as a controllable $\mathcal{PT}$ symmetry. Then, we perform a multi-method study of this model, focusing on the low-energy spectrum, the thermodynamic quantities, and the transport properties associated with different fixed points. Using the Bethe ansatz approach, we identify the existence of the Yu-Shiba-Rusinov-like state previously found in the non-Hermitian single-channel Kondo model. Then, based on non-Hermitian numerical renormalization group calculations, we reveal clear numerical signatures of the Yu-Shiba-Rusinov state emerging in the relatively strong non-Hermiticity regime of the $\mathcal{PT}$-asymmetric model. Furthermore, our boundary conformal field theory, which is found to be applicable for the $\mathcal{PT}$-symmetric model, uncovers an anomalous temperature dependence of the Kondo conductance, which is beyond conventional Hermitian Kondo systems.
@@ I Introduction
Non-Hermitian quantum models, which physically arise from the loss or gain via coupling to an external environment, are well known to host a wealth of quantum phenomena absent in the Hermitian counterparts, e.g., the non Hermitian skin effect [1, 2, 3, 4, 5, 6, 7, 8, 9], non-Hermitian topology [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], exceptional points [20, 21, 22, 23, 24] and unexpected dynamics [25, 26, 27, 28, 29, 30, 31]. These findings open a new avenue in which many key concepts or tools for closed systems find new applications. However, most existing works investigate non-Hermitian effects in non-interacting or weakly interacting systems, where quasi-particle descriptions are well-defined and applicable. Although there has been some studies on correlated non-Hermitian systems, they mainly focus on specific solvable models [32, 33, 34, 35, 36, 37] as well as their ground state properties. The question of how the non-Hermiticity would affect correlated systems without well-defined particles has not been addressed yet. Particularly, it remains unclear whether or not any universal collective behavior would emerge in such cases.
@@
A typical and important example that violates the quasi-particle description is the non-Fermi liquid (NFL). It is one of the core concepts in modern condensed matter physics, describing a novel class of emergent phenomena beyond the Landau Fermi liquid paradigm. In strong-correlated systems where the conventional quasi-particle picture breaks down, there could emerge exotic collective behaviors, such as the unconventional scaling of thermodynamic quantities versus temperatures. A thorough understanding of these correlated phenomena is of key importance, as they are fundamentally related to the high-temperature superconductors [38, 39, 40, 41, 42, 43, 44], quantum spin liquids [45, 46, 47, 48, 49, 50], topological orders [51, 52, 53, 54], etc. On the other hand, the mechanism driving the non-Fermi liquid behavior can be various. It could arise from strong coupling between electrons and critical bosons, thus could naturally take place around quantum criticalities [55, 56, 57]. Besides, it can also be driven by the proximity to quantum spin liquids [58, 59, 60, 61, 62, 63], which indicates that the frustration and correlation are essential for its generation.
@@
Experimental signatures of NFL have been observed in various correlated systems, including heavy fermion materials [64, 65, 66, 67, 68, 69, 70, 71, 72, 73], high-$T_{c}$ superconductors [38, 39, 40, 41, 42, 43, 44], etc. Among them, the multichannel Kondo effect has attracted enormous interest, as it serves as prototypical platform for the experimental control over the NFL behaviors [74, 75, 76, 77, 78, 79]. To realize a multichannel Kondo system that drives the NFL fixed point, exact channel symmetry should be achieved. Although this has been experimentally demonstrated by fine-tuning in nanostructures [74, 75], another remarkable proposal is to utilize the topological Kondo effect. Based on multi-junction setup where Majorana fermions tunnel into normal leads, Béri and Cooper demonstrate a feasible and controllable way to realize the multi-channel Kondo model with exact channel symmetry without the need of fine-tuning [80]. Considering that the multichannel Kondo model hosts non-Fermi liquid fixed point, which exhibits fractionalized degrees of freedom contributing to a fractional impurity entropy [81, 82, 83], it is intriguing and worthwhile to ask would any new collective physics be emerging when non-Hermiticity further comes into play?
@@
Despite the multichannel feature, the non-Hermitian single channel Kondo problem has been proposed and studied in previous works  [35, 37]. Ref. [37] carried out a perturbative renormalization group calculations and found that there occur two fixed points, respectively in the weak- and strong-coupling regime. Due to the non-Hermitian effects, there occurs a circling RG flow around the weak-coupling fixed point, driving the system towards the local moment phase. Moreover, it is argued that the strong-coupling regime describes the Kondo phase, in analogy with conventional Kondo physics of the Hermitian counterpart. A more recent study [35] then performed a more careful solution of the Bethe ansatz solution. Remarkably, it reveals a more interesting phase, where the impurity is screened by effective extra degrees of freedom arising from the non-Hermitian effect, forming a single-mode bound state. Owing to its similarity with the Yu-Shiba-Rusinov (YSR) impurity states in conventional $s$-wave superconductors, it is dubbed the YSR phase. These results indicate that the non-Hermiticity brings about a richer phase diagram than expected, which is worth a detailed and comprehensive study by non-perturbative methods.
@@
In this work, we perform a comprehensive, multi-method investigation of the non-Hermitian Kondo model, with a particular focus on the multichannel cases. Our aim here is threefold. First, although the multi-junction setup [80] can be used to realize the conventional Hermitian multichannel Kondo model, it remains unclear how to systematically introduce the non-Hermitian effects into the system. Thus, we propose a realistic setup and demonstrate a step-by-step realization scheme. Second, we use a number of methods to thoroughly study the non-Hermitian multichannel Kondo (NHMCK) model in order to achieve a more comprehensive understanding beyond perturbative perspectives, involving the perturbative renormalization group (RG), Bethe ansatz, non-Hermitian numerical renormalization group (NHNRG), non-Hermitian Kubo formula, as well as the boundary conformal field theory (BCFT). Particularly, we clarify the role of $\mathcal{PT}$ symmetry by considering both the $\mathcal{PT}$-symmetric and $\mathcal{PT}$-asymmetric NHMCK. Different features in terms of the low-energy spectrum and the thermodynamic quantities are found for the two cases. For the $\mathcal{PT}$-asymmetric NHMCK, using the NHNRG method, originally proposed by Ref.[84], we reveal numerical evidenve for the YSR phase predicted by Ref. [35]. For the $\mathcal{PT}$-symmetric model, we combine the NHNRG with an additional analytic proof and show that the low-energy physics in this case is well within the BCFT descriptions. Third, we go beyond previous work and explore the nontrivial transport behaviors generated by the NFL and the non-Hermitian physics. We find that both the weak-coupling and the strong-coupling fixed point exhibit unconventional Kondo conductance different from the Hermitian systems. Around the weak-coupling FP, the impurity moment decouples from the conduction electrons even for antiferromagnetic Kondo couplings, leading to a universal low-temperature conductance, $\sim 1/\mathrm{ln}^{2}(T/T_{K})$. Although the conductance is similar to that of the underscreened Kondo models [85, 86, 87], it is driven by a completely different dissipation mechanism–the nH-induced decoupling. Remarkably, around the strong-coupling fixed point, the BCFT analysis predicts an anomalous deviation in the Kondo conductance with raising temperature. Such a non-Hermiticity enriched anomaly stems from an emergent BCFT with non-Hermitian boundary operators, which is unique to systems exhibiting both non-Hermiticity and non-Fermi liquid behaviors.
@@
Last, we mention that this present work is a companion to another work by us, i.e., Ref. [88]. More concise results with more clear pictures are discussed in Ref. [88], while the technical aspects in terms of methods and more detailed discussions are included in this manuscript.
II summary of results
@@
In this section, we will outline the content in the remaining part of the work and our main results. First, we will introduce the non-Hermitian multichannel Kondo Hamiltonian and demonstrate a step-by-step realization scheme of this model, as presented in Sec. III. Then, in Sec. IV, we perform a perturbative RG analysis, which suggests two possible phases corresponding to the weak- and strong-coupling fixed points. They are referred to as the weak-coupling local moment phase and the strong coupling Kondo phase , respectively. In Sec. V, we present the Bethe ansatz calculations that reveal the existence of an intermediate YSR-like phase, which is beyond the perturbative RG analysis. Then, to better confirm these phases, we carry out a NHNRG study of the model. In Sec. VI, we first introduce the technical details of the implementation of the NHNRG method and then utilize it to investigate the $\mathcal{PT}$-asymmetric NHMCK model. Our focus in Sec. VII is on the pseudo-Hermitian model with $\mathcal{PT}$ symmetry. We first show the NHNRG results for the low-energy spectrum and the thermodynamic quantities, which justify the validity of the strong-coupling Kondo phase predicted by the perturbative RG. Then, we outline a more rigorous proof showing that the BCFT approach still applies for this model with $\mathcal{PT}$ symmetry. In Sec. VIII, we explore the transport properties associated with both fixed points. For the weak-coupling fixed point, we present the calculation procedures based on the non-Hermitian Kubo formula, which is well applicable for the perturbative regime. Then, for the strong-coupling fixed point, we exploit the BCFT, and demonstrate the anomalous Kondo conductance generated by the interplay between non-Hermiticity and non-Fermi liquid physics. More discussions and conclusions are included in Sec. IX.
     weak-coupling intermediate strong-coupling
local-moment phase YSR-like phase Kondo phase
[…]
Bethe ansatz        $\checkmark$       $\checkmark$      $\checkmark$
NHNRG        $\checkmark,\ {\color[rgb]{1,0,0}\bm{T}}$       $\checkmark,\ {\color[rgb]{1,0,0}\bm{T}}$      $\checkmark,\ \color[rgb]{1,0,0}\bm{T}$
[…]
For clarity, we now summarize our main results, focusing on the three possible phases, i.e., the weak-coupling local moment phase, the intermediate YSR-like phase, and the strong coupling Kondo phase, and how they could be revealed by different analytical and numerical approaches.
First, the perturbative RG analysis up to 2-loop order reveals two fixed points, which are the local-moment fixed point and the strong-coupling Kondo fixed point, as mentioned above. Based on the RG flow of Kondo coupling, we further adopt the non-Hermitian Kubo formula to calculate the conductance around the weak-coupling fixed point. However, we note that the perturbative approach cannot yield useful information about the intermediate YSR-like phase. Meanwhile, the predictions about the strong-coupling fixed point needs to be further confirmed since it is essentially beyond the perturbation theory.
Second, by means of rigorously solving the Bethe ansatz equations of the non-Hermitian multichannel Kondo model, we identify all the three distinct phases, controlled by a non-Hermiticity parameter $\theta$. The intermediate YSR-like phase is characterized by a single-mode bound state that contributes to the impurity screening, which is beyond the reach of the perturbative RG analysis.
Third, using the NHNRG method, we further calculate the thermodynamic properties in both the UV and IR regimes with varying $\theta$. From the obtained results including the impurity entropy and the specific heat, we not only arrive at the weak-coupling and strong-coupling phases, but also uncover numerical signatures of the YSR-like bound state. These results are in support of the Bethe ansatz solutions, indicating a richer phase diagram beyond the scope of Hermitian Kondo models.
Fourth, we analyze the quantum transport properties in the weak-coupling local-moment phase and the strong-coupling Kondo phase for both the UV and IR limits. The conductance in the weak-coupling regime is obtained perturbatively from a non-Hermitian (NH) extension of the Kubo formula. Since this NH Kubo formula cannot be applied to non-perturbative regimes, we restrict ourselves to $\mathcal{PT}$-symmetric models and utilize the non-perturbative BCFT techniques to extract the IR behavior of the conductance in the strong-coupling Kondo phase. This approach also yields thermodynamic quantities for the strong-coupling Kondo phase, consistent with the results obtained by the NHNRG approach.
All the above results are concisely demonstrated by Table. 1, which contains a compact “cross-validation” summary of the three phases from a multi-method perspective.
III Model Hamiltonian and possible realization setup
@@
Figure 1: The schematic plot of the realization setup of the NHMCK model. (a) The multi-junction nanostructure, where the Majorana modes (the red dot) on a superconducting island are connected to normal leads. The tunneling between the Majorana mode and the lead is assisted by a quantum dot (the orange dot). (b) A zoom-in plot of the Majorana-lead tunneling junction via the quantum dot. The quantum dot is further coupled to an environment. (c) The quantum dot has a single-body loss into the environment and monitored by a reservoir, and a post-selection is implemented on the Lindbladian dynamics of the quantum dot.
The tunneling junction is plotted in more details in Fig.1(b), which involves the dissipative QD. The latter consists of a QD coupled to environment. The QD-environment coupling system is further subjected to a monitoring-feedback device, which resets the particle loss events and ensures the particle conservation of the QD, as shown by Fig.1(c).
@@
where $\varrho$ is the reduced density matrix of the QD, $H^{\prime}=H_{\mathrm{QD}}+H_{\mathrm{LS}}$ where $H_{\mathrm{LS}}$ is the Lamb-shift term taking the form of $H_{\mathrm{LS}}=\sum_{i}\Omega_{i}d_{i}^{\dagger}d_{i}$, with $\Omega_{i}$ being a pesudo-potential for QD acquired from coupling to the environment. The jump operators $L_{i}$, which originate from the coupling of the quantum dot operators to the bath, can be constructed through the following manner, [93, 94, 91, 95],
@@
In order to realize a non-Hermitian system, the particle leakage event should be monitored by an attached empty reservoir as a particle detector [90, 97] and then an immediate resetting needs to be implemented by the feedback equipment. An alternative option is to take the record in the time period when no particle leakage is monitored. These procedures result in a postselection on the Lindbladian, leading to an effective non-Hermitian Hamiltonian description of the QD, i.e.,
@@
where $\tilde{\epsilon}_{d,i}$ denotes the Lamb-shifted QD energy level. It is clear from above that the coupling to the environment and the postselection generate an dissipative QD with an imaginary chemical potential. Note that the effective non-Hermitian Hamiltonian description can be also accessed through the Green’s function point of view (see Appendix A).
@@
This normalized density matrix will be used later in our calculation of Kondo conductance, as will be shown in Sec. VIII.
@@
where we have focused on the single-mode case with $N_{d}=1$ thus neglected the index $i$ for simplicity. The Majorana-lead tunneling is given by,
@@
We now show that the dissipative effect of the QD can induce a non-reciprocal tunneling process of the Majorana fermions. This can be read off from the Schrödinger equation of the junction system, i.e., $H_{\mathrm{junc}}|\Psi\rangle=E|\Psi\rangle$, which is written as a coupled equation in the direct product space of the QD and the rest part. Tracing out the QD degrees of freedom then generates an effective junction Hamiltonian renormalized by the dissipative QD as,
@@
Inserting of Eq.(8), we find that the original tunneling amplitude is renormalized in a nontrivial way. The tunneling from the Majorana to the lead ($t$) and the reversed tunneling $(t^{\star})$ are respectively modified as (see Appendix B),
@@
IV Perturbative renormalization group results